joe is 9 years younger than maria. in 4 years, she will be twice as old as he will be. find their percent age?
Answer by Maths68(1474)


Let
Present age of Joe = j
Present age of Maria = m
joe is 9 years younger than maria.
m=j+9.......(1)
In 4 years,
m+4=j+4
then
she will be twice as old as he will be
m+4=2(j+4)
m+4=2j+8
m=2j+8-4
m=2j+4
Put the value of m from (1) to above equation
j+9=2j+4
j-2j=4-9
-j=-5
Multiply by (-1) both sides
(-1)(-j)=(-1)(-5)
j=5
Put the value of j in (1)
m=j+9.......(1)
m=5+9
m=14
Present age of Joe = j = 5 years old
Present age of Maria = m = 14 years old